# Baking Gaussian Splatting into Diffusion Denoiser for Fast and Scalable Single-stage Image-to-3D Generation and Reconstruction

### v3

@@ Abstract
Existing feedforward image-to-3D methods mainly rely on 2D multi-view diffusion models that cannot guarantee 3D consistency. These methods easily collapse when changing the prompt view direction and mainly handle object-centric cases. In this paper, we propose a novel single-stage 3D diffusion model, DiffusionGS, for object generation and scene reconstruction from a single view. DiffusionGS directly outputs 3D Gaussian point clouds at each timestep to enforce view consistency and allow the model to generate robustly given prompt views of any directions, beyond object-centric inputs. Plus, to improve the capability and generality of DiffusionGS, we scale up 3D training data by developing a scene-object mixed training strategy. Experiments show that DiffusionGS yields improvements of 2.20 dB/23.25 and 1.34 dB/19.16 in PSNR/FID for objects and scenes than the state-of-the-art methods, without depth estimator. Plus, our method enjoys over 5$\times$ faster speed ($\sim$6s on an A100 GPU). Project page: https://caiyuanhao1998.github.io/project/DiffusionGS/
@@ 3.1 DiffusionGS
$\small\mathbf{x}_{t}^{(i)}=\bar{\alpha_{t}}\mathbf{x}_{0}^{(i)}+\sqrt{1-\bar{\alpha_{t}}}\epsilon_{t}^{(i)},\vspace{-1.5mm}$ (1)
@@
$\small\mathcal{G}_{\theta}(\mathcal{X}_{t}|\mathbf{x}_{con},\mathbf{v}_{con},t,\mathcal{V})=\{G_{t}^{(k)}(\bm{\mu}_{t}^{(k)},\mathbf{\Sigma}_{t}^{(k)},\alpha_{t}^{(k)},\bm{c}_{t}^{(k)})\},\vspace{-0.5mm}$ (2)
@@
$\small u_{t}^{(k)}=w_{t}^{(k)}u_{near}+(1-w_{t}^{(k)})u_{far},\vspace{-1mm}$ (3)
@@
Denoiser Architecture. As shown in Fig. 4 (b), the input images concatenated with the viewpoint conditions are patchified, linearly projected, and then concatenated with a positional embedding to derive the input tokens of the Transformer backbone, which consists of $L$ blocks. Each block contains a multi-head self-attention (MSA), an MLP, and two layer normalization (LN). Eventually, the output tokens are fed into the Gaussian decoder to be linearly projected and then unpatchified into per-pixel Gaussian maps ^=$\hat{\mathcal{H}}=\{\hat{\mathbf{H}}_{con},\hat{\mathbf{H}}^{(1)},\cdots,\hat{\mathbf{H}}^{(N)}\}$, where $\hat{\mathbf{H}}_{con}$ and $\hat{\mathbf{H}}^{(i)}\in\mathbb{R}^{H\times W\times 14}$. Then $N+1$ Gaussian maps are merged into the Gaussian point clouds $\mathcal{G}_{\theta}$ in Eq. (2). The timestep condition controls the Transformer block and Gaussian decorder through the adaptive layer normalization mechanism [52].
@@
$\small\hat{\mathbf{x}}_{(0,t)}^{(i)}=F_{r}(\mathbf{M}_{ext}^{(i)},\mathbf{M}_{int}^{(i)},\mathcal{G}_{\theta}(\mathcal{X}_{t}|\mathbf{x}_{con},\mathbf{v}_{con},t,\mathcal{V})),\vspace{-1mm}$ (4)
@@
$\small\mathbf{\bf\Sigma^{\prime}}_{t}^{(k,i)}=\mathbf{J}_{t}^{(i)}\mathbf{W}_{t}^{(i)}\mathbf{\Sigma}_{t}^{(k)}{\mathbf{W}_{t}^{(i)}}^{\top}{\mathbf{J}_{t}^{(i)}}^{\top},\vspace{-1.8mm}$ (5)
@@
$\small\mathcal{L}_{de}=\mathcal{L}_{2}(\hat{\mathcal{X}}_{(0,t)},{\mathcal{X}}_{0})+\lambda\cdot\mathcal{L}_{\text{VGG}}(\hat{\mathcal{X}}_{(0,t)},{\mathcal{X}}_{0}).\vspace{-0.8mm}$ (7)
In the testing phase, our DiffusionGS randomly samples noise from standard normal distribution at timestep $T$ and then gradually denoise it step by step. The predicted $\hat{\mathcal{X}}_{(0,t)}$ at each timestep $t$ is fed into the next timestep $t-1$ to replace $\mathcal{X}_{0}$ in Eq. (1) for sampling $\mathcal{X}_{t-1}$ at each noisy view as
$\small\mathbf{x}_{t-1}^{(i)}=\bar{\alpha}_{t-1}\mathbf{\hat{x}}_{(0,t)}^{(i)}+\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{t-1}^{(i)}.\vspace{0mm}$ (8)
@@ 3.2 Scene-Object Mixed Training Strategy
$\small\theta_{cd}^{(i)}\leq\theta_{1},~{}~{}\theta_{dn}^{(i,j)}\leq\theta_{2},\vspace{-1.5mm}$ (9)
@@
$\small\frac{\vec{z}_{con}\cdot\vec{z}_{noise}^{~{}(i)}}{|\vec{z}_{con}|\cdot|\vec{z}_{noise}^{~{}(i)}|}\geq\text{cos}(\varphi_{1}),~{}~{}\frac{\vec{z}_{con}\cdot\vec{z}_{nv}^{~{}(j)}}{|\vec{z}_{con}|\cdot|\vec{z}_{nv}^{~{}(j)}|}\geq\text{cos}(\varphi_{2}),\vspace{-0.6mm}$ (10)
@@
$\small\bm{r}=(\bm{o}-(\bm{o}\cdot\bm{d})\bm{d},\bm{d})\vspace{-1.2mm}$ (11)
@@
Dual Gaussian Decoder. As the depth range varies across object- and scene-level datasets, we use two MLPs to decode the Gaussian primitives for objects and scenes in mixed training. As shown in Fig. 4 (b), for the object-level Gaussian decoder, the nearest and farthest distances [$u_{near}$, $u_{far}$] in Eq. (3) are set as [0.1, 4.2] and $\bm{\mu}_{t}^{(k)}$ is clipped into $[-1,1]^{3}$. For the scene-level Gaussian decoder, [$u_{near}$, $u_{far}$] is set to [0, 500]. The two decoders are also controlled by the timestep embedding. In the finetuning phase, we just use a single decoder while the other is removed.
@@
[…]
@@
$\small\mathcal{L}=(\mathcal{L}_{de}+\mathcal{L}_{nv})\cdot\mathbf{1}_{\text{iter}>\text{iter}_{0}}+\mathcal{L}_{pd}\cdot\mathbf{1}_{\text{iter}\leq\text{iter}_{0}}\cdot\mathbf{1}_{\text{object}},\vspace{-1.2mm}$ (13)
@@ 4 Experiment
Dataset. We use Objaverse [12] and MVImgNet [82] as the training sets for objects. We center and scale each 3D object of Objaverse into $[-1,1]^{3}$, and render 32 images at random viewpoints with 50 FOV. For MVImgNet, we crop the object, remove the background, normalize the cameras, and center and scale the object to $[-1,1]^{3}$. We preprocess 730K and 220K training samples in Objaverse and MVImgNet. We use the ABO [11] and GSO [13] datasets for evaluation. We adopt RealEstate10K [90] and DL3DV10K [34] as the scene-level training datasets. RealEstate10K includes 80K video clips of indoor and outdoor real scenes selected from YouTube videos. We follow the standard training/testing split. DL3DV10K contains 10510 videos of real-world scenarios, covering 96 complex categories. For all evaluation, each instance has 1 input view and 10 testing views.
Implementation Details. We implement DiffusionGS by Pytorch [51] and train it with Adam optimizer [27]. To save GPU memory, we adopt mixed-precision training [48] with BF16, sublinear memory training [8], and deferred GS rendering [86]. In mixed training, we use 32 A100 GPUs to train the model on Objaverse, MVImgNet, RealEstate10K, and DL3DV10K for 40K iterations at the per-GPU batch size of 16. Then we finetune the model on the object- and scene-level datasets with 64 A100 GPUs for 80K and 54K iterations at the per-GPU batch size of 8 and 16. The learning rate is linearly warmed up to 4e-4 with 2K iterations and decays to 0 using cosine annealing scheme [42]. Finally, we scale up the training resolution from 256$\times$256 to 512$\times$512 and finetune the model for 20K iterations.
@@ 4.1 Comparison with State-of-the-art Methods
We conduct a user study by inviting 25 people to score the visual quality of the generation results of 14 objects according to the 3D geometry, texture quality, and alignment with the prompt view. The user study score ranges from 1 (worst) to 6 (best). Tab. 7(a) reports the results and running time at the size of 256$\times$256. Our method achieves the highest score while enjoying over 5$\times$ and 10$\times$ inference speed compared to the recent best 3D diffusion DMV3D and multi-view diffusion-based method 12345++. Tab. 7(c) and 7(d) list the results of object generation on the ABO and GSO datasets. DiffusionGS surpasses DMV3D by 2.2/1.25 dB in PSNR and 23.25/21.96 in FID score on ABO/GSO.

### v5

@@ Abstract
[…]
@@ 3.1 DiffusionGS
$\small\mathbf{x}_{t}^{(i)}=\bar{\alpha_{t}}\mathbf{x}_{0}^{(i)}+\sqrt{1-\bar{\alpha_{t}}}\epsilon_{t}^{(i)},\vskip-4.2679pt$ (1)
@@
$\small\mathcal{G}_{\theta}(\mathcal{X}_{t}|\mathbf{x}_{con},\mathbf{v}_{con},t,\mathcal{V})=\{G_{t}^{(k)}(\bm{\mu}_{t}^{(k)},\mathbf{\Sigma}_{t}^{(k)},\alpha_{t}^{(k)},\bm{c}_{t}^{(k)})\},\vskip-1.42262pt$ (2)
@@
$\small u_{t}^{(k)}=w_{t}^{(k)}u_{near}+(1-w_{t}^{(k)})u_{far},\vskip-2.84526pt$ (3)
@@
Denoiser Architecture. As shown in Fig. 4 (b), the input images concatenated with the viewpoint conditions are patchified, linearly projected, and then concatenated with a positional embedding to derive the input tokens of the Transformer backbone, which consists of $L$ blocks. Each block contains a multi-head self-attention (MSA), an MLP, and two layer normalization (LN). Eventually, the output tokens are fed into the Gaussian decoder to be linearly projected and then unpatchified into per-pixel Gaussian maps $\hat{\mathcal{H}}=\{\hat{\mathbf{H}}_{con},\hat{\mathbf{H}}^{(1)},\cdots,\hat{\mathbf{H}}^{(N)}\}$, where $\hat{\mathbf{H}}_{con}$ and $\hat{\mathbf{H}}^{(i)}\in\mathbb{R}^{H\times W\times 14}$. Then $N+1$ Gaussian maps are merged into the Gaussian point clouds $\mathcal{G}_{\theta}$ in Eq. (2). The timestep condition controls the Transformer block and Gaussian decorder through the adaptive layer normalization mechanism [52].
@@
$\small\hat{\mathbf{x}}_{(0,t)}^{(i)}=F_{r}(\mathbf{M}_{ext}^{(i)},\mathbf{M}_{int}^{(i)},\mathcal{G}_{\theta}(\mathcal{X}_{t}|\mathbf{x}_{con},\mathbf{v}_{con},t,\mathcal{V})),\vskip-2.84526pt$ (4)
@@
$\small\mathbf{\bf\Sigma^{\prime}}_{t}^{(k,i)}=\mathbf{J}_{t}^{(i)}\mathbf{W}_{t}^{(i)}\mathbf{\Sigma}_{t}^{(k)}{\mathbf{W}_{t}^{(i)}}^{\top}{\mathbf{J}_{t}^{(i)}}^{\top},\vskip-5.12149pt$ (5)
@@
$\small\mathcal{L}_{de}=\mathcal{L}_{2}(\hat{\mathcal{X}}_{(0,t)},{\mathcal{X}}_{0})+\lambda\cdot\mathcal{L}_{\text{VGG}}(\hat{\mathcal{X}}_{(0,t)},{\mathcal{X}}_{0}).\vskip-2.27621pt$ (7)
In the testing phase, our DiffusionGS randomly samples noise from standard normal distribution at timestep $T$ and then gradually denoise it step by step. The predicted $\hat{\mathcal{X}}_{(0,t)}$ at each timestep $t$ is fed into the next timestep t1t-1 to replace $\mathcal{X}_{0}$ in Eq. (1) for sampling $\mathcal{X}_{t-1}$ at each noisy view as
$\small\mathbf{x}_{t-1}^{(i)}=\bar{\alpha}_{t-1}\mathbf{\hat{x}}_{(0,t)}^{(i)}+\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{t-1}^{(i)}.\vskip 0.0pt$ (8)
@@ 3.2 Scene-Object Mixed Training Strategy
$\small\theta_{cd}^{(i)}\leq\theta_{1},~~\theta_{dn}^{(i,j)}\leq\theta_{2},\vskip-4.2679pt$ (9)
@@
$\small\frac{\vec{z}_{con}\cdot\vec{z}_{noise}^{~(i)}}{|\vec{z}_{con}|\cdot|\vec{z}_{noise}^{~(i)}|}\geq\text{cos}(\varphi_{1}),~~\frac{\vec{z}_{con}\cdot\vec{z}_{nv}^{~(j)}}{|\vec{z}_{con}|\cdot|\vec{z}_{nv}^{~(j)}|}\geq\text{cos}(\varphi_{2}),\vskip-1.70717pt$ (10)
@@
$\small\bm{r}=(\bm{o}-(\bm{o}\cdot\bm{d})\bm{d},\bm{d})\vskip-3.41432pt$ (11)
@@
Dual Gaussian Decoder. As the depth range varies across object- and scene-level datasets, we use two MLPs to decode the Gaussian primitives for objects and scenes in mixed training. As shown in Fig. 4 (b), for the object-level Gaussian decoder, the nearest and farthest distances $u_{near}$, $u_{far}$] in Eq. (3) are set as [0.1, 4.2] and $\bm{\mu}_{t}^{(k)}$ is clipped into $1,1]^{3}$. For the scene-level Gaussian decoder, $u_{near}$, $u_{far}$] is set to [0, 500]. The two decoders are also controlled by the timestep embedding. In the finetuning phase, we just use a single decoder while the other is removed.
@@
$\small\mathcal{L}_{pd}=\underset{k}{\mathbb{E}}[l_{t}^{(k)}-(\frac{l_{t}^{(k)}-\underset{k}{\mathbb{E}}[l_{t}^{(k)}]}{\sqrt{\text{Var}(l_{t}^{(k)})}}\sigma_{0}+\underset{k}{\mathbb{E}}[|\bm{o}^{(k)}|])],\vskip-5.69054pt$ (12)
@@
$\small\mathcal{L}=(\mathcal{L}_{de}+\mathcal{L}_{nv})\cdot\mathbf{1}_{\text{iter}>\text{iter}_{0}}+\mathcal{L}_{pd}\cdot\mathbf{1}_{\text{iter}\leq\text{iter}_{0}}\cdot\mathbf{1}_{\text{object}},\vskip-3.41432pt$ (13)
@@ 4 Experiment
Dataset. We use Objaverse [12] and MVImgNet [82] as the training sets for objects. We center and scale each 3D object of Objaverse into $1,1]^{3}$, and render 32 images at random viewpoints with 50 FOV. For MVImgNet, we crop the object, remove the background, normalize the cameras, and center and scale the object to $1,1]^{3}$. We preprocess 730K and 220K training samples in Objaverse and MVImgNet. We use the ABO [11] and GSO [13] datasets for evaluation. We adopt RealEstate10K [90] and DL3DV10K [34] as the scene-level training datasets. RealEstate10K includes 80K video clips of indoor and outdoor real scenes selected from YouTube videos. We follow the standard training/testing split. DL3DV10K contains 10510 videos of real-world scenarios, covering 96 complex categories. For all evaluation, each instance has 1 input view and 10 testing views.
Implementation Details. We implement DiffusionGS by Pytorch [51] and train it with Adam optimizer [27]. To save GPU memory, we adopt mixed-precision training [48] with BF16, sublinear memory training [8], and deferred GS rendering [86]. In mixed training, we use 32 A100 GPUs to train the model on Objaverse, MVImgNet, RealEstate10K, and DL3DV10K for 40K iterations at the per-GPU batch size of 16. Then we finetune the model on the object- and scene-level datasets with 64 A100 GPUs for 80K and 54K iterations at the per-GPU batch size of 8 and 16. The learning rate is linearly warmed up to 4e-4 with 2K iterations and decays to 0 using cosine annealing scheme [42]. Finally, we scale up the training resolution from $\times256$ to $\times512$ and finetune the model for 20K iterations.
@@ 4.1 Comparison with State-of-the-art Methods
[…]